A man throws ball to maximum horizontal distance of 80m. Calculate maximum height reached..?

Grade:11

1 Answers

Arun
6 years ago
We know that R = u2 * Sin 2θ / g.
As, the range is maximum, so, θ = 450
So, 80 = u2 /g.
Now, H = height reached by ball =  u2 Sin2θ / 2g.
Putting value of θ = 45, u2 / g = 80, we get
H = 80/4  = 20m